OpenStudy (anonymous):

x+y=0 3x-3y=0 (-1,1)

OpenStudy (anonymous):

yes that is correct because when u plug in the point into the function you get 0=0

OpenStudy (anonymous):

i get -1=0,-6=0

OpenStudy (anonymous):

no you get -1+1=0 when you plug in the coordinates to the first function and for the second you get 3(-1)-3(1)=0 you get 0=0.

OpenStudy (anonymous):

there both true because left side =right side.

OpenStudy (anonymous):

i thought 3(-1) =-3

OpenStudy (anonymous):

this what i did 3(-1)-3(1)=0 -3-3=0 -6=0

OpenStudy (anonymous):

sorry @jinan you right i forgot about the other negative for the second function. So the top function is true and the one below it is false.

OpenStudy (anonymous):

x+y=0 this what i get for this one -1+1=0 -1=0 (-1,1)

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

no that one is not correct because they cancel each other out. -1+1 =0

OpenStudy (anonymous):

-1+1=0 0=0

OpenStudy (anonymous):

ok you are right that you so much

OpenStudy (anonymous):

your welcome! @jinan

OpenStudy (anonymous):

do you know anything about graphing

OpenStudy (anonymous):

i have two questions for graphing

OpenStudy (anonymous):

what are the question?

OpenStudy (anonymous):

-x+y=1 x+y=0

OpenStudy (anonymous):

graph the linear

OpenStudy (anonymous):

sorry that wasnt right

OpenStudy (anonymous):

-x+x=-2x right????????

OpenStudy (anonymous):

y+y=2y

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

hod on i'm trying to solve it right now. give me a minute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u have an idea on how you graph is suppose to look?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

 does that look familiar